OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x – 5y + z = 12 x + 6y – z = –4 2x – 5y – 2z = –8 4x + y = 8 –17y = 0 2x – 11y = 16 4x + 7y = –16 –23y + 4z = 24 –17y = 0 13y – 2z = 24 7y – 4z = 0

OpenStudy (anonymous):

@rajathsbhat

OpenStudy (anonymous):

what you have to do is re-arrange the second equation to get x (in terms of y & z). Once you've done that, substitute that 'x' into wherever there is x in the first & third equation. Simplify to get 2 equations in y & z. Try it out!

OpenStudy (anonymous):

x=-6y+z-4 so then... 3(-6y+z-4)-5y+z=12 -18y+3z-12-5y+z=12 -23y+4z=24 and... 2x–5y–2z=–8 2(-6y+z-4)-5y-2z=-8 -12y+2z-8-5y-2z=-8 -17y=0

OpenStudy (anonymous):

C! :)

OpenStudy (anonymous):

thanks @rajathsbhat

OpenStudy (anonymous):

sure thing, yummy :)